OpenStudy (anonymous):

How do you prove 1=0 using integration by parts?

OpenStudy (kainui):

You can do it, you just gotta start with this integral \[\Large \int\limits \frac{1}{x}dx\]

OpenStudy (anonymous):

its very serious situation :D

OpenStudy (kainui):

\[\Large x^{-1}=u \ \ dx=dv \\ \Large -x^{-2}dx=du \ \ x=v\]

OpenStudy (anonymous):

oh then what xD

OpenStudy (kainui):

\[\Large \int\limits \frac{1}{x}dx=uv-\int\limits vdu \\ \Large \frac{1}{x}x-\int\limits x \frac{-1}{x^2}dx \\ \Large 1+ \int\limits \frac{1}{x} dx\]

OpenStudy (anonymous):

i see

OpenStudy (kainui):

Now all you gotta do is just see that they are equal, and subtract off the integrals. So it's obvious now that all of calculus is just a lie

OpenStudy (anonymous):

infact thats also proof 0=1,2,3,4,5,......n for any integer xD

OpenStudy (kainui):

Well true, but once you realize that 1=0 then we can apply this to itself: \[0=1=1+0=1+1=2 = 2+0 = 2+1 = 3 =...\]

OpenStudy (kainui):

Pretty handy so now we can go ahead and prove that 0 = infinty

OpenStudy (anonymous):

yep interesting !

OpenStudy (kainui):

Luckily though it's not a big deal, since \[n=n*1=n*0=0\] So it makes sense, since all numbers are really zero all along!

OpenStudy (kainui):

This appeals to our common sense, I've never "seen" a 3, I've seen 3 books but just because I can hold a book doesn't mean I can hold a 3. So obviously numbers don't exist by this proof and it makes sense since no one's ever held a number in their hands before, it's pretty obvious.

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

seems like all illusion xD they made the whole thing to troll kids in school by studying unnecessary stuff bahahaha

OpenStudy (anonymous):

now lets see why this integral invalid :D

OpenStudy (anonymous):

do you know why ?

OpenStudy (kainui):

Yeahh foolish kids! Yeah I know why it's easy once you know what the trick is XD

OpenStudy (anonymous):

i know xD sure u know

OpenStudy (kainui):

The indefinite integral of the copy function is CTRL + C

OpenStudy (anonymous):

@Abhisar @amandanewsomm @bohotness @Brad1996 @bazel_e @BriWhtThtMouthDo_0722 @Compassionate @Coolsector @CausticSyndicalist @Donblue22 @DarkMoonZ @damonte212 @DawnStar53 @eliassaab @eric_d @Elsa213 @fallenangelorchid @ganeshie8 @ganeshie8 @hba @haleyelizabeth2017 @ispike @iambatman @Joel_the_boss @jojo4eva @just_one_last_goodbye @Jonnychewy @jademc01 @Kainui @KamiBug @kaylala @kristyn-whitaker @Kittymeowb @KimJongUn @leonardo0430 @mathmate @micahm @marylou004 @morganchavez @mcnabkt @Marki @Nnesha @neko_of_hearts @Nicksl99 @One098 @Palmo4ka @Radelp @rishavraj @sumanta @Shadowgirl15 @scburchell11 @soccer130130 @Sherabartlett @sciencewiz3000 @TuringTest @Teddyiswatshecallsme @TwiztTiez @TFIOS_GUS623 @TheRaggedyDoctor @TheSmartOne @uri @uChezzy @Werewolfprincess

OpenStudy (bohotness):

yes

OpenStudy (one098):

What's with it and the mass tagging, honestly?

OpenStudy (leonardo0430):

too much tagging m8 ,-,

OpenStudy (bohotness):

its against coc and thats spamming tags

OpenStudy (fallenangelorchid):

The bloody hell with the mass tagging.

OpenStudy (donblue22):

Please do not mass tag it is against the CoC

OpenStudy (one098):

lol Logan XD

OpenStudy (fallenangelorchid):

xDDD

OpenStudy (anonymous):

so any help from u nice awesome folks :D

Nnesha (nnesha):

@Marki i got you ":) very intelligent haan :

OpenStudy (anonymous):

Can people please stop putting me in mass tags? I have work I need help with on here and I don't need notifications distracting me. Sorry if I came across rude. thanks.

OpenStudy (anonymous):

just help me :'( ans say why 0=1 in this question

OpenStudy (bohotness):

sure

ganeshie8 (ganeshie8):

I like this fake proof xD we can prove 0 equals any number using this trick : \[\int f(x) dx \stackrel{?}{=} \int f(x) dx + C \] "subtract" and be happy :P \[0 \stackrel{?}{=} C\]

OpenStudy (anonymous):

:D yes indeed

OpenStudy (kainui):

Yeah but that ruins the fun of it haha. The whole point is to trick students who are in calculus 2 into thinking integration by parts is mysterious when really it's just the product rule lol

OpenStudy (anonymous):

.